Can you show me the steps on how to do this equation?
x/5 = (x-2)/3

Answer by kngt(2)

Your question is x/5=(x-2)/3
Here x/5 is the Left Hand Side (LHS) of your Equation
and (x-2)/3 is the Right Hand Side (RHS) of your Equation.
When 5 goes to the RHS '/5' is converted to '*5' ie, division is converted to multiplication.
Similarly, when 3 goes to the LHS '/3' is converted to '*3'.(Similarly '+' is converted to '-' and '-' is converted to '+')
This method is called cross multiplication.
So the new equation is
3*x = 5*(x-2)
3*x = 5*x - 5*2 (opening the bracket)
3*x - 5*x = -10 (5*x goes to LHS so + sign is converted to - sign)
-2*x = -10
x = -10/-2 ( a negative no. divided by another negative no. is a positive no.)
x = 5